OpenStudy (anonymous):

Identify the correct slope and y-intercept: of the equation 2x + 3y = 6 pls help

OpenStudy (jdoe0001):

solve for "y", that is isolate "y" and see what you'd get, then compare it agains the "slope-intercept" form => http://www.algebra-class.com/image-files/slope-formula-1.gif

OpenStudy (anonymous):

slope = -2; y-intercept: = (0, 6) slope= -3/2; y-intercept = (0,3) slope = -2; y-intercept: = (0, 3) slope = -2/3; y-intercept = (0,2)

OpenStudy (anonymous):

those are the options

OpenStudy (anonymous):

it has to be on of those four answers